### v2

# LaTeXTrans: Structured LaTeX Translation with Multi-Agent Coordination

Abstract
Despite the remarkable progress of modern machine translation (MT) systems on general-domain texts, translating structured LaTeX-formatted documents remains a significant challenge. These documents typically interleave natural language with domain-specific syntax, such as mathematical equations, tables, figures, and cross-references, all of which must be accurately preserved to maintain semantic integrity and compilability. In this paper, we introduce LaTeXTrans, a collaborative multi-agent system designed to address this challenge. LaTeXTrans ensures format preservation, structural fidelity, and terminology consistency through six specialized agents: 1) a Parser that decomposes LaTeX into translation-friendly units via placeholder substitution and syntax filtering; 2) a Translator, Validator, Summarizer, and Terminology Extractor that work collaboratively to ensure context-aware, self-correcting, and terminology-consistent translations; 3) a Generator that reconstructs the translated content into well-structured LaTeX documents. Experimental results demonstrate that LaTeXTrans can outperform mainstream MT systems in both translation accuracy and structural fidelity, offering an effective and practical solution for translating LaTeX-formatted documents. The code of LaTeXTrans is available at https://github.com/NiuTrans/LaTeXTrans.
LaTeXTrans: Structured LaTeX Translation
with Multi-Agent Coordination
Ziming Zhu1††thanks: Authors contributed equally., Chenglong Wang1∗, Shunjie Xing1, Yifu Huo1, Fengning Tian2, Quan Du2, Di Yang1,2, Chunliang Zhang1,2, Tong Xiao1,2††thanks: Corresponding author. and Jingbo Zhu1,2 1 School of Computer Science and Engineering, Northeastern University, Shenyang, China 2 NiuTrans Research, Shenyang, China {zhuzm0721, clwang1119}@gmail.com, {xiaotong, zhujingbo}@mail.neu.edu.cn
1 Introduction
LaTeX is a widely adopted macro package system built on top of TeX, designed to facilitate the typesetting of complex and structured documents. It has become the de facto standard for scholarly publications across a wide range of scientific disciplines. According to recent statistics, nearly 98% of scientific papers are published in English, while only about 3% of the global population speaks English as their first language (Kleidermacher and Zou, 2025). This linguistic disparity places considerable pressure on non-native English speakers, who are frequently required to read or write LaTeX-formatted documents in English. As a result, the technical barriers to academic learning and research are significantly increased.
A straightforward approach to ease this burden is to translate LaTeX documents into the user’s native language by processing the compiled PDF version, a process referred to as PDF translation. However, this approach often results in incomplete formatting due to errors in PDF parsing. In contrast, a more promising alternative is to translate directly at the LaTeX source level and then compile the translated content into a target-language PDF document. This approach can preserve structural information and allows better control over formatting.
However, translating LaTeX source files presents unique challenges not encountered in plain-text translation. LaTeX documents interleave natural language with domain-specific markup, such as mathematical equations, citation commands, and formatting environments, all of which must be precisely preserved to ensure semantic correctness and successful compilation. Naively applying standard MT systems to LaTeX code typically leads to broken syntax, semantic errors, or formatting loss, ultimately hindering rather than helping the user.
To address these challenges, in this paper, we introduce LaTeXTrans, a collaborative multi-agent system designed to directly translate LaTeX source files while preserving their structural and semantic integrity. Our LaTeXTrans operates on raw LaTeX code and maintains the full syntactic and semantic structure of the document throughout the entire translation pipeline. Specifically, it comprises three modules and six specialized agents:
Parsing Module: Responsible for fine-grained analysis of LaTeX-formatted documents. To handle the structural complexity of LaTeX, we design a Parser agent equipped with a placeholder mechanism and a syntax filter, which together decompose the source into manageable translation units.
Translation Module: This module leverages a team of collaborative agents, including a Translator, Validator, Summarizer, and Terminology Extractor, which work together to perform context-aware and self-correcting translation of the parsed units.
Generation Module: A Generator agent reconstructs the translated document by reinserting the translated content into the original LaTeX structure, producing well-formatted LaTeX source in the target language.
To evaluate the effectiveness of LaTeXTrans, we first construct a LaTeX source test set using TeX files collected from arXiv papers. We then compare LaTeXTrans with a range of MT and LLM-based translation baselines. Experimental results demonstrate that LaTeXTrans consistently outperforms all baselines in both translation accuracy and format fidelity. Notably, LaTeXTrans achieves an improvement of 13.20 points on FC-score, along with significant gains in COMETkiwi and LLM-score when compared to GPT-4o.
2 Related works
LLM-based Machine Translation.
The emergence of LLMs has introduced a new paradigm for MT, shifting away from traditional supervised learning on parallel corpora toward more flexible, general-purpose language understanding (Gain et al., 2025). LLMs like GPT-3 (Brown et al., 2020), PaLM (Chowdhery et al., 2022), and GPT-4 demonstrate strong multilingual capabilities without explicit training on translation tasks. LLM-based translation leverages in-context learning, where the model is prompted with examples or instructions to perform translation on the fly. This approach has shown competitive performance in zero-shot and few-shot learning scenarios (Vilar et al., 2023; Luo et al., 2025), especially for high-resource language pairs. Unlike traditional neural machine translation (NMT), which requires retraining or fine-tuning for each new domain or language, LLMs can generalize across tasks and languages with minimal additional data.
Multi-Agent Systems.
More recently, the emergence of LLMs has opened new possibilities for Multi-Agent Systems (MAS). In LLM-based multi-agent systems, each agent is instantiated as an LLM-powered entity capable of natural language reasoning, planning, and collaboration. Systems such as AutoGPT (Yang et al., 2023), CAMEL (Li et al., 2023), and AutoGen (Dibia et al., 2024) demonstrate that LLM agents can simulate diverse roles and complete complex tasks through dialogue-based coordination. A growing number of studies explore the use of multi-agent systems for translation-related tasks. Notably, MAS has emerged as a promising solution for document-level translation (Wang et al., 2024), a long-standing challenge in MT.
Formatted Text Translation.
Formatted text translation involves translating documents that contain structural or semantic markup, such as LaTeX and XML. These formats often interleave natural language with commands, tags, or tokens that encode formatting, layout, or semantic annotations. Although some recent efforts have been made in this direction Kleidermacher and Zou (2025); Khan (2025), formatted text translation still faces two major challenges. The first is the lack of a robust, general-purpose system specifically designed for translating formatted content. Currently, only a few proprietary tools, such as Youdao and Baidu, offer relatively effective solutions. While open-source tools like MathTranslate***https://github.com/SUSYUSTC/MathTranslate and GPT-Academic†††https://github.com/binary-husky/gpt_academic have received positive feedback, they still lag behind commercial systems in overall performance. The second is the lack of a sound, formatted text translation evaluation technique. As traditional BLEU or COMET scores do not cover format correctness or tag retention. Therefore, it is imperative to develop a new evaluation technique for structure-aware translation.
3 System Design
The key architecture of LaTeXTrans is a multi-agent coordination designed for translating structured LaTeX documents. It consists of three modules: the Parser, the Translation Module, and the Generation Module. The design and functionality of each component are described in detail below.
3.1 Parser Module
Structured LaTeX documents interleave natural language content with formatting commands and semantic markup, resulting in tightly coupled representations that are not well-suited for direct translation by LLMs. Naively feeding the entire document to an LLM leads to several issues: unnecessary processing of non-translatable components, increased computational cost, and a higher risk of introducing translation errors. To address these challenges, we introduce the Parser module, which serves as the first stage of the LaTeXTrans pipeline. Its basic idea is to transform complex LaTeX documents into clean, structured translation units that are easier for LLMs to process. Specifically, we design a placeholder substitution strategy to temporarily replace LaTeX-specific commands and environments, and implement a filtering mechanism to remove components that do not require translation.
Placeholder Substitution Strategy.
For a common LaTeX document, our placeholder substitution strategy is shown in Figure 2. We consider that the original mathematical formulas and charts are retained during translation. The first step is to replace the captions in the chart with placeholders. The second step is to replace the environment with placeholders, which will include the vast majority of mathematical formulas, charts, and other parts that do not need to be translated. Finally, we split the replaced text into sections (including subsections and subsubsections). For a LaTeX project composed of multiple tex files, we first merge the necessary tex files into the main file and then insert placeholders at the beginning and end of the merge for future restoration. The subsequent placeholder replacement rules and segmentation methods are the same as before. From the placeholder substitution strategy, we obtain translation units of two granularities: context (i.e., section and environment) and sentence (i.e., caption).
Translation Unit Filter.
While non-translatable components are replaced with placeholders, we notice that LaTeX allows users to define custom environments, making it infeasible to rely solely on exhaustive rule-based approaches to identify all such segments. To address this issue, we complement a predefined list of protected environments with a Filter agent powered by an LLM, which dynamically determines whether a given environment requires translation. Each extracted environment is annotated with a binary label: True or False. The translation module subsequently processes only those segments labeled as True.
3.2 Translation Module
The translation module comprises four agents: the Translator, Validator, Summarizer, and Terminology Extractor. After the Translator completes the translation of all designated units, the output is passed to the Validator, which generates an error report and returns it for revision if necessary. The Summarizer and Terminology Extractor assist the Translator by providing a summary of the preceding content and a domain-specific terminology dictionary, respectively, thereby enhancing contextual coherence and ensuring terminology consistency throughout the translation process.
Translator-Validator Iteration.
[…]
Summarizer and Terminology Extractor.
Inspired by Wang et al. (2024)’s work, we design a Summarizer and Terminology Extractor to enhance the contextual coherence and terminology consistency of translation. Specifically, the Summarizer is responsible for constantly generating and updating the summary of the previous text during the translation process. When each translation unit is completed, the Summarizer will combine the previous summary with the original text of the current translation unit to generate a new summary. The Terminology Extractor is responsible for maintaining a terminology dictionary and adding it to the prompt of the Translator to provide a reference for terminology translation for the Translator. When the Translator finishes the translation of a translation unit, the Terminology Extractor extracts term pairs from the original text and the translation and updates the term dictionary in real-time.
3.3 Generation Module
The generation module is responsible for reassembling the translation units into structured LaTeX documents and compiling the structured LaTeX documents into PDF files using specific compilers (e.g. pdfLaTeX and XeLaTeX).
4 Experiment
| System | En-Zh | En-Ja | ||||||
| Cometkiwi () | LLM-score () | FC-score () | Cost () | Cometkiwi () | LLM-score () | FC-score () | Cost () | |
| NiuTrans | 64.69 | 7.93 | 60.72 | - | 65.49 | 8.19 | 27.48 | - |
| Google Translate | 46.23 | 5.93 | 51.00 | - | 56.21 | 7.01 | 50.00 | - |
| LLaMA-3.1-8b | 42.89 | 2.92 | 49.40 | - | 44.49 | 3.32 | 60.92 | - |
| Qwen-3-8b | 45.55 | 7.87 | 48.68 | - | 46.20 | 6.80 | 49.52 | - |
| Qwen-3-14b | 68.18 | 8.76 | 65.63 | - | 72.84 | 8.66 | 61.88 | - |
| DeepSeek-V3 | 67.26 | 9.02 | 63.68 | $0.02 | 72.17 | 9.00 | 63.96 | $0.03 |
| GPT-4o | 67.22 | 8.58 | 58.32 | $0.13 | 71.16 | 8.91 | 56.92 | $0.11 |
| LaTeXTransQwen-3-14b | 71.37 | 8.97 | 71.20 | - | 74.68 | 8.51 | 59.84 | - |
| LaTeXTransDeepSeek-V3 | 73.48 | 9.01 | 70.52 | $0.10 | 75.39 | 8.89 | 66.52 | $0.13 |
| LaTeXTransGPT-4o | 73.59 | 8.92 | 71.52 | $0.35 | 74.47 | 8.93 | 64.92 | $0.45 |
4.1 Settings
Datasets.
Since no publicly available LaTeX document dataset currently exists, we constructed our test set by selecting the TeX sources of 50 English academic papers from the arXiv repository. The chosen papers include both long and short articles, many of which contain complex formulas and figures, ensuring structural diversity and complexity in the LaTeX content. Further experimental details are provided in Appendix A.
Baselines.
Our baselines are categorized into two groups: traditional MT systems and LLM-based translation systems. For the former, we selected NiuTrans and Google Translate as representative systems. For the latter, we evaluated five strong LLMs, including both open-source and proprietary models: LLaMA-3.1-8B Grattafiori et al. (2024), Qwen-3-8B Yang et al. (2025), Qwen-3-14B, DeepSeek-V3 Liu et al. (2024), and GPT-4o Hurst et al. (2024). Among these, Qwen-3-14B, DeepSeek-V3, and GPT-4o were further used as the backbone models for agents in LaTeXTrans.
4.2 Evaluation Metrics
We conducted a comprehensive assessment of our system from two dimensions: translation quality and format retention ability.
Translation Quality.
Because high-quality reference translations for LaTeX documents require expert-level annotation, we adopted wmt22-cometkiwi-da Rei et al. (2022), a reference-free evaluation metric (denoted as Cometkiwi), to assess the translation quality of LaTeX documents. Furthermore, we employed GPT-4o as an automatic evaluator to further assess translation quality across multiple dimensions, guided by carefully designed system prompts. The evaluation covered four aspects: Faithfulness, Fluency, Terminology Consistency, and Coherence, where each was rated on a scale from 0 to 10. An overall score was then synthesized by GPT-4o based on the individual scores across these dimensions (denoted as LLM-score).
Format Retention Ability.
Whether the labels are completely retained is an important manifestation of the ability of the formatted text translation system. However, at present, there is no universal indicator to evaluate the format retention ability of models or systems during the translation process. Therefore, for LaTeX documents, we have designed a new evaluation metric, Format Consistency Score (denoted as FC-score), to assess the retention ability of our system for LaTeX labels during the translation process. We can compute the FC-score by
where is the initial score before the rewards and penalties, is the penalty coefficient per error, is the penalty coefficient per warning, is the reward for successful compilation. and are numbers of errors and warnings, indicates whether the LaTeX document compiled successfully. We then clip the score to the valid range , and are the upper bound and lower bound of the score (e.g. 0~100).
4.3 Results
We evaluate our LaTeXTrans system on two translation tasks: English-to-Chinese (En-Zh) and English-to-Japanese (En-Ja). The results, shown in Table 1, demonstrate that LaTeXTrans consistently outperforms both the NMT and Single-Agent baselines across all evaluation metrics, including COMETkiwi and FC-score. In terms of translation quality, LaTeXTrans demonstrates substantial improvements in FC-score (71.52 vs. 58.32 for the En-Zh task and 70.52 vs. 63.68 for the En-Ja task), indicating significantly better preservation of LaTeX formatting during translation. Moreover, when powered by GPT-4o as the backbone model, LaTeXTrans achieves the highest scores across all three evaluation metrics—COMETkiwi, LLM-score, and FC-score—underscoring its strong overall translation performance on structured LaTeX documents. In terms of translation cost, LaTeXTrans delivers superior performance without incurring a substantial increase in computational expense compared to other LLM-based translation systems, making it well-suited for large-scale deployment in real-world applications.
4.4 Ablation Study
Table 2 presents an ablation study on the En–Zh task using GPT-4o and DeepSeek-V3 as backbone models. Introducing the Parser module significantly improves both COMETkiwi and FC-score, indicating that the placeholder substitution strategy enhances translation quality and label preservation. Adding the Validator module further boosts overall performance, although a slight drop in LLM-score is observed with DeepSeek-V3. We hypothesize that this is due to the Validator enforcing strict tag retention through iterative checks, which may restrict the Translator and slightly impact fluency. Finally, incorporating the Summarizer and Terminology Extractor improves the LLM-score, reflecting better cross-paragraph coherence. However, slight declines in COMETkiwi and FC-score suggest that these improvements may not be fully captured by COMETkiwi. A detailed analysis with a case study is provided in Section 4.4.1.
| Setting | GPT-4o | DeepSeek-V3 | ||||
| Cometkiwi | LLM-score | FC-score | Cometkiwi | LLM-score | FC-score | |
| SA. (Baseline) | 67.22 | 8.58 | 58.32 | 67.26 | 9.02 | 63.68 |
| SA. + P. | 74.47 | 8.89 | 69.64 | 74.39 | 9.03 | 70.08 |
| SA. + P. + V. | 74.57 | 8.91 | 71.76 | 74.42 | 8.94 | 70.80 |
| SA. + P. + V. + S. | 74.06 | 8.95 | 71.64 | 74.02 | 9.05 | 70.68 |
| SA. + P. + V. + S. + TE. | 73.59 | 8.93 | 71.52 | 73.48 | 9.01 | 70.52 |
4.4.1 Translation consistency
We present a case study of the En-Zh task from our test set to demonstrate that our system does indeed perform better in terms of translation consistency, as shown in Figure 3. In this case, the terminology translation of LaTeXTrans remains consistent across the three sections. In contrast, the baseline method finds it difficult to maintain such consistency. This indicates that our system can maintain excellent consistency throughout the entire translation process.
5 Conclusion
In this paper, we propose LaTeXTrans, a multi-agent system for translating structured LaTeX documents. LaTeXTrans consists of three collaborative modules, each responsible for a specific stage of the translation pipeline. Experimental results demonstrate that LaTeXTrans can outperform baseline systems and offer a reliable solution for LaTeX document translation.
Limitations
Any instruction-following LLM can be integrated into our LaTeXTrans system. However, due to the large number of available models, it is impractical to evaluate each one individually. Therefore, we select a representative subset of commonly used LLMs for our experiments. We believe this selection sufficiently demonstrates the practicality and effectiveness of LaTeXTrans for LaTeX document translation. Additionally, although commercial systems such as Baidu and Youdao offer LaTeX translation services, they are not open-source. As a result, we are unable to compute metrics like COMETkiwi and FC-score for these systems. Therefore, we do not include a comprehensive comparison with them in our main experiments.
Appendix A Additional Detailed Settings of the Experiment
Baselines.
Since the dataset consisted entirely of structured LaTeX documents which exceeded the handling capabilities of single-model systems, we adopted a preprocessing step in the baseline approach. Specifically, the structured LaTeX documents were segmented into section-level translation units to make them manageable for translation.
Hyperparameter Setting.
In the experiments, we evaluated both open-source and closed-source models separately. For the closed-source models, we accessed them via a third-party API. In the baseline approach, we set the maximum number of new tokens to 16,384 and the temperature to 0.7, while keeping all other hyperparameters at their default values. For our system, the temperature in the Filter was set to 0 with a maximum of 50 new tokens, while all other agents were configured with a maximum of 8,192 new tokens; the remaining hyperparameters were kept at their defaults.
Evaluation.
When computing COMETkiwi and LLM-scores, we used pylatexenc‡‡‡https://github.com/phfaist/pylatexenc to convert each LaTeX translation unit into plain text. Although LaTeXTrans parses structured LaTeX documents into fine-grained translation units, we followed the baseline’s evaluation protocol by using section-level translation units for computing both COMETkiwi and LLM-scores. Furthermore, to assess contextual consistency in the LLM-score evaluation, we concatenated section-level translation units into paired paragraphs and then scored them using GPT-4o. The prompt template used for scoring is illustrated in Figure 12. When calculating the FC-score, we set the initial score to 100. Since errors have a greater impact on the final PDF format scheduling effect than warnings, in the experiment, we set the value of (10) to be significantly greater than (2). Ultimately, whether the compilation is successful is the most intuitive factor for evaluating the compilation. Therefore, in the experiment, we set the to 20.
Datasets
We selected the LaTeX source files of 50 academic papers in the field of computer science from arXiv as our test set. The distribution of paper lengths is shown in Figure 4. Additionally, we analyzed the topics of the papers and visualized them as a word cloud in Figure 5. This result shows that the test set exhibits a diverse range of paper lengths, covering both short and long documents, which helps ensure robustness across different document sizes. Moreover, the word cloud reveals a wide variety of research topics within the computer science domain, confirming the topical diversity of the test set and enhancing the generality of our evaluation.
Appendix B System Performance Display
We select six cases to visually demonstrate the translation performance of our system, focusing on En-Zh and En-Ja translation tasks, as illustrated in Figure 6 to Figure 11. All six cases are translation cases of the LaTeX source code of papers by LaTeXTrans. In each case, we have selected two relatively complex parts to present. Among the six cases, there are three En-Zh translation tasks and three En-Ja translation tasks, respectively.
Appendix C Prompt Templates for LLM-Based Components in LaTeXTrans